# Brightest Cluster Galaxy Offsets in Cold Dark Matter

### v1

@@ 5.1 Dependence on Softening Length
In Fig. 2 we present these offset distributions across all simulations considered, as measured by SE centroids. Here it can be seen that the majority of the offset distributions are below the softening length scale in each case, and that the IllustrisTNG 300-3 and BAHAMAS distributions are consistent, while possessing almost identical softening lengths. There also exists a clear trend of decreasing offsets with decreasing softening length, which is the expectation in a physics model whose “true” offset distribution is consistent with zero, or at least smaller than the smallest softening length considered here, namely $1.48\,\rm{kpc}$ in IllustrisTNG 300-1. We also represent the median and 68th percentile offsets as a function of softening length in the inset panel of Fig. 2. Here it is made clear that with centroids measured by SourceExtractor, the offset measurements are limited by the gravitational softening and convergence properties, consistent with the findings of Schaller et al. (2015b), Ng et al. (2017) and Harvey et al. (2019) in which offsets measured between the DM centroid (or potential minimum) and BCG are typically on the order of or below the softening length. The lower resolution simulations do exhibit some offsets above softening, but these may be due to resolution and zoom box nuisance parameters, as the relatively poor mass resolution necessitates a larger zoom box used to create the mass maps. As discussed in Sec. 5.5 the size of the zoom box positively correlates with the median offset, as measuring the center on a larger scale has the potential to deviate due to substructure.
@@ 5.2 Dependence on Centroid Identification
Finally, the large scale gas-BCG offsets are the largest of any centroid identification method, indicating that the gas is a poor tracer of the stellar, and therefore dark matter in the core. This can also be seen in Fig. 1, as in the large-scale gas the centroid is offset from the potential minimum by $\sim 50\,\rm{kpc}$, while the small scale structure of the gas is highly clumpy, making a small scale measurement of the gas centroid only loosely correlated with the underlying potential. That the gas is an unreliable tracer of the potential minimum in simulations is a robust statement as the offsets measured are above the convergence radius of the underlying dark matter. Though it is possible that the gas physics is improperly resolved on small scales, the large scale gas centroid is more relevant to observations and is a poorer tracer of the potential minimum than even the small scale centroids. Thus inferences made using the gas as a tracer for the potential center, on both large and small scales, can lead to misinterpretations of the underlying dark matter physics.
@@ 5.3 Dependence on Redshift and Mass Scale
Figure 5: Density profiles of example halos in full-physics IllustrisTNG300 runs at three resolution levels, with the density profiles of their matched dark matter (DM)-only run counterparts. Shown are the total density profiles and DM density profiles in the full-physics run, and the DM density profiles in the DM-only run. Halos are also matched across simulations (along columns) in both mass and radius to below displayed precision. The convergence radius is calculated as $2.8\epsilon_{\rm{DM}}$ where $\epsilon_{\rm{DM}}$ is the gravitational softening length (see text).
@@ 5.4 Dependence on Baryons
In Fig. 5 it can be seen that the effect of the baryonic model on the DM density profiles (dotted lines) is to increase the density in the core, consistent with the adiabatic contraction caused by gas condensation and mass accretion seen in other simulations (eg. Schaller et al., 2015a), but only below the convergence radius. Since these changes take place primarily below the convergence radius, robust claims about the impact of this baryonic model on offset measurements cannot be made. For example offsets are overwhelmingly measured below the softening scale, which may be in part due to increased central densities as a result of this contraction, but this cannot be concluded reliably from this data. It is also the case that no significant trend appears in cluster size, with the evidence of contraction being comparable at all mass levels shown.
@@
Figure 6: Nuisance parameter tests for BCG offset measurements presented as cumulative probabilities, all performed in IllustrisTNG300-1 with the fiducial parameters of Sec. 3 and Sec. 4 unless otherwise specified. The cumulative probability is the probability that an offset is measured as less than or equal to the corresponding $x$-axis value. All offsets calculated via SourceExtractor centroids for stars and dark matter, except for the lensing scale panel. Note the difference in scale for the rightmost column. (Upper left) depth of the zoom box used to create the mass maps along the line of sight. (Upper middle) pixel size used to generate mass maps. Note that for this test only groups above $3\times 10^{14}\,\rm{M_{\odot}}$ were considered, justified by the lack of dependence of the distribution on mass. (Upper right) side length of the zoom box used to create the mass maps. (Lower left) sightline used for creation of mass maps, for example “$z$” means the mass map was calculated from the perspective of the positive $z$-axis. (Lower middle) threshold used in SourceExtractor detection algorithm, in units of the standard deviation of the identified background. (Lower right) Einstein radius used as a lensing scale for the area-matched isophote method. Lensing scale tests are all performed with a $300\,\rm{kpc}$ field of view. Arrows on a panel represent the value of the parameter for that offset distribution (matched in color) where appropriate.
@@ 5.5 Parameter Tests
    Pixel Size: We test pixel sizes both significantly smaller and larger than the choice adopted in this paper, but find no dependence of the offset distribution on this choice above the softening length. Below the softening length there is a non-monotonic variance in the distribution, but this regime is unconverged and thus will not affect inference made with these results. In the case of $10\,\rm{kpc}$ pixels the source extraction and thresholding are likely breaking down, with the background window sizes and filters approaching the size of the mass map.
@@ 6 Conclusions
To constrain the nature of dark matter, and in particular its self-interaction cross section, it is necessary to improve the simulation data, analysis techniques and the catalogue of observations. In simulations, improvements can be made by examining high time resolution data to study the dynamics in the cores of clusters at low gravitational softening length. In the analysis of the data, it will be valuable to conduct extensive mock lensing observations in simulated clusters to establish the CDM BCG-strong lensing offset distribution. Finally, the catalogue of observations with high-quality lens models for which the BCG position is not chosen to coincide with the DM halo center has great potential to be expanded both in the near term via a homogeneous re-analysis of Hubble archival data, and in the future via large lensing surveys conducted by the Euclid and Roman space telescopes.
Acknowledgements
This work makes use of the following software: Python (Van Rossum & Drake, 2009), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Astropy Collaboration et al., 2013, 2018), jupyter (Kluyver et al., 2016). This research also made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2023). CR thanks Meredith Neyer for numerous helpful discussions.
SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1].

### v3

@@ 5.1 Dependence on Softening Length
In Fig. 2 we present these offset distributions across all simulations considered, as measured by SE centroids. Here it can be seen that the majority of the offset distributions are below the softening length scale in each case, and that the IllustrisTNG 300-3 and BAHAMAS distributions are consistent, while possessing almost identical softening lengths. There also exists a clear trend of decreasing offsets with decreasing softening length, which is the expectation in a physics model whose “true” offset distribution is consistent with zero, or at least smaller than the smallest softening length considered here, namely $1.48\,\rm{kpc}$ in IllustrisTNG 300-1. We also represent the median and 68th percentile offsets as a function of softening length in the inset panel of Fig. 2. Here it is made clear that with centroids measured by SourceExtractor, the offset measurements are limited by the gravitational softening and convergence properties, consistent with the findings of Schaller et al. (2015a), Ng et al. (2017) and Harvey et al. (2019) in which offsets measured between the DM centroid (or potential minimum) and BCG are typically on the order of or below the softening length.
The lower resolution simulations do exhibit some offsets above softening, but these may be due to resolution and zoom box nuisance parameters, as the relatively poor mass resolution necessitates a larger zoom box used to create the mass maps. As discussed in Sec. 5.5 the size of the zoom box positively correlates with the median offset, as measuring the center on a larger scale has the potential to deviate due to substructure. A small amount of offsets above softening are observed in all distributions, and can often be identified by visual inspection as caused by merging cores and bimodal mass distributions, or a failure of the source extraction to faithfully represent the large scale halo in rare cases. We leave a detailed study of the physical nature of large offsets to future work, focusing instead here on the limitations of CDM simulations and the rarity of offsets above softening.
@@ 5.2 Dependence on Centroid Identification
Finally, the large scale gas-BCG offsets are the largest of any centroid identification method, indicating that the gas is a poor tracer of the stellar, and therefore dark matter in the core. This can also be seen in Fig. 1, as in the large-scale gas the centroid is offset from the potential minimum by $\sim 50\,\rm{kpc}$, while the small scale structure of the gas is highly clumpy, making a small scale measurement of the gas centroid only loosely correlated with the underlying potential. These findings are qualitatively in line with those of Seppi et al. (2023), in which the BCG-gas offsets were measured by various means in simulations (including TNG300-1) and in observations, as a means of estimating cluster disturbance. The BCG was defined in Seppi et al. (2023) to lie at the potential minimum, which we demonstrate to be coincident above the softening scale with the BCG in Fig. 3, and the gas center was computed via an emission measure weighted center of mass. The gas-BCG offsets in this paper are however systematically smaller by a factor of $\sim 2$, likely owing to the significant difference in centroid identification methodology. The scale of these offsets is also consistent with observations such as those of Lauer et al. (2014) in which 433 gas-BCG offsets were observed, 15% of which were above $100\,\rm{kpc}$. The offsets observed here are again slightly smaller than those observed at 15% above $\sim 60\,\rm{kpc}$, but it should be noted that many observational effects are not modelled here.
That the gas is an unreliable tracer of the potential minimum in simulations is a robust statement as the offsets measured are above the convergence radius of the underlying dark matter. Though it is possible that the gas physics is improperly resolved on small scales, the large scale gas centroid is more relevant to observations and is a poorer tracer of the potential minimum than even the small scale centroids. Thus inferences made using the gas as a tracer for the potential center, on both large and small scales, can lead to misinterpretations of the underlying dark matter physics.
@@ 5.3 Dependence on Redshift and Mass Scale
Figure 5: Density profiles of example halos in full-physics IllustrisTNG300 runs at three resolution levels, with the density profiles of their matched dark matter (DM)-only run counterparts. Shown are the total density profiles and DM density profiles in the full-physics run, and the DM density profiles in the DM-only run. Halos are also matched across simulations (along columns) in both mass and radius to below displayed precision, to demonstrate how similar halos look at each resolution level. Halo IDs in the group catalogs of the full-physics and DM-only runs are also provided. The convergence radius is calculated as $2.8\epsilon_{\rm{DM}}$ where $\epsilon_{\rm{DM}}$ is the gravitational softening length (see text).
@@ 5.4 Dependence on Baryons
In Fig. 5 it can be seen that the effect of the baryonic model on the DM density profiles (dotted lines) is to increase the density in the core, consistent with the adiabatic contraction caused by gas condensation and mass accretion seen in other simulations (eg. Schaller et al., 2015b), but only below the convergence radius. Since these changes take place primarily below the convergence radius, robust claims about the impact of this baryonic model on offset measurements cannot be made. For example offsets are overwhelmingly measured below the softening scale, which may be in part due to increased central densities as a result of this contraction, but this cannot be concluded reliably from this data. It is also the case that no significant trend appears in cluster size, with the evidence of contraction being comparable at all mass levels shown.
@@
Figure 6: Nuisance parameter tests for BCG offset measurements presented as cumulative probabilities, all performed in IllustrisTNG300-1 with the fiducial parameters of Sec. 3 and Sec. 4 unless otherwise specified. The cumulative probability is the probability that an offset is measured as less than or equal to the corresponding $x$-axis value. All offsets calculated via SourceExtractor centroids for stars and dark matter, except for the lensing scale panel. Note the difference in scale for the rightmost column. (Upper left) depth of the zoom box used to create the mass maps along the line of sight. (Upper middle) pixel size used to generate mass maps. (Upper right) side length of the zoom box used to create the mass maps. (Lower left) sightline used for creation of mass maps, for example “$z$” means the mass map was calculated from the perspective of the positive $z$-axis. (Lower middle) threshold used in SourceExtractor detection algorithm, in units of the standard deviation of the identified background. (Lower right) Einstein radius used as a lensing scale for the area-matched isophote method. Lensing scale tests are all performed with a $300\,\rm{kpc}$ field of view. Arrows on a panel represent the value of the parameter for that offset distribution (matched in color) where appropriate.
@@ 5.5 Parameter Tests
    Pixel Size: We test pixel sizes up to five times smaller and larger than the choice adopted in this paper ($0.25\,\rm{kpc}$ on each side), but find no dependence of the offset distribution on this choice above the softening length. Below the softening length there is a non-monotonic variance in the distribution, but this regime is unconverged and thus will not affect inference made with these results.
@@ 6 Conclusions
To constrain the nature of dark matter, and in particular its self-interaction cross section, it is necessary to improve the simulation data, analysis techniques and the catalogue of observations. In simulations, improvements can be made by examining high time resolution data to study the dynamics in the cores of clusters at low gravitational softening length. In the analysis of the data, it will be valuable to conduct extensive mock lensing observations in simulated clusters to establish the CDM BCG-strong lensing offset distribution. Finally, the catalogue of observations with high-quality lens models for which the BCG position is not chosen to coincide with the DM halo center has great potential to be expanded: in the near term via a homogeneous re-analysis of Hubble archival data and highly-constrained James Webb space telescope images of strong lensing systems (eg. Cha et al. (2024)), and in the future via large lensing surveys conducted by the Euclid and Nancy Grace Roman space telescopes.
Acknowledgments
CR thanks Meredith Neyer for numerous helpful discussions. SB is supported by the UK Research and Innovation (UKRI) Future Leaders Fellowship [grant number MR/V023381/1].
This work makes use of the following software: Python (Van Rossum & Drake, 2009), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), astropy (Astropy Collaboration et al., 2013, 2018), jupyter (Kluyver et al., 2016). This research also made use of Photutils, an Astropy package for detection and photometry of astronomical sources (Bradley et al., 2023).